shyhoneyyy:

A rocket is released a jet fighter flying horizontally at 1200 kph at an altitude of 2400m above its target. The rocket thrust gives it a constant acceleration of 0.6g. Determine the angle between the horizontal and the line of the sight to the target.

Narad:

Initial velocity of rocket is = u=1200*1000/3600= 333.3m/s Horizontal acceleration of rocket is a=0.6 m/s^2 Horizontal distance to the target is d=333.3*t+0.5*0.6g*t^2 Vertical acceleration of rocket is = g m/s^2 Initial vertical velocity of rocket is u =0m/s Therefore, 2400=0*t+0.5*g*t^2 t^2=2400/(0.5*g)=489.8s^2 t=sqrt(489.8)=22.13s Therefore, d=333.3*22.13+0.5*0.6*g*22.13^2 = 8816.4 m The angle is = arctan(2400/8816.4)=15.2ยบ